FAR-INFRARED LASER STARK SPECTROSCOPY OF ^{13}CD_{3}OD

\underline{M. JACKSON}, Department of Physics, New Mexico State University, Las Cruces, NM 88003; G. R. SUDHAKARAN, E. GANSEN, Department of Physics, University of Wisconsin, La Crosse, WI 54601; I. MUKHOPADHYAY, Laser Programme, Centre for Advanced Technology, Indore, India, 452 013; and R. M. LEES, Department of Physics, University of New Brunswick, Fredericton, New Brunswick, Canada E3B 5A3.

The Stark spectrum of ^{13}CD_{3}OD has been investigated in the far-infrared region using the HCN and DCN lasers. The spectra were taken at room temperature for both parallel and perpendicular polarizations using electric fields up to 60 000 \frac{V}{cm}. An extensive series of absorption lines observed with the 337 μm line has been assigned to the J = 9 to 15 members of the K = 5 ← 4 E_{2} Q-branch in the υ_{t} = 1 torsional state^{a}. Further analysis of this Q-branch multiplet has yielded the branch origin, ν_{0} = 882 847.7 MHz, and expansion coefficients, a = 93.6785 MHz and b = −0.2689 MHz. An R-branch transition observed at 337 μm has been assigned as J_{K} = 9_{7} ← 8_{6} A υ_{t} = 0. Two families of resonances observed with the 195 μm line have been tentatively identified as members of the Q-branch multiplet J_{K} = J_{7} ← J_{8} E_{2} υ_{t} = 2 for J = 11 and 12. Zero-field frequencies for all assigned transitions are given with improved accuracy over those calculated from the available molecular constants.

aM. Jackson, G. R. Sudhakaran, E. Gansen, J. Mol. Spectrosc., 176, 439-441 (1996).